# PETSc/TAO Developments for Early Exascale Systems

### v1

@@ 4 Communication on GPUs
As PETSc’s strength lies in distributed computation with sparse and irregular data, the library has to take care of complex communication among processes, such as those in sparse matrix-vector multiplication, sparse matrix-matrix multiplication, or irregular mesh partitioning. In PETSc, we designed a module named PetscSF to encapsulate frequently used communication operations. Underneath a unified interface, PetscSF can have different communication implementations, though the default uses MPI. In this section, we first introduce PetscSF’s design and the extensions to allow MPI communication with device data. Next we introduce an experimental PetscSF implementation using NVIDIA NVSHMEM (NVIDIA, 2024) that can overcome some limitations imposed by MPI. Finally, we use a microbenchmark written with PetscSF to study the communication performance on the four target machines.
@@ 4.1 The star-forest (SF) abstraction
PetscSF uses star-forests to abstract communication patterns. A star is a simple tree consisting of one root vertex connected to zero or more leaves. A star forest is a disjoint union of stars, see examples in Figure 1.
@@
The former broadcasts root values to leaves, while the latter reduces leaf values into roots, with both taking a reduction operation specified by an MPI_Op argument to add the source values or to replace the destination values (with op=MPI_REPLACE). The Begin/End split-phase design allows users to insert computations in between to potentially overlap computation with communication. It is common in irregular applications that leaves and roots are not consecutive in their respective index space. In that case, PETSc will call its pack or unpack kernels to put data into internally managed buffers for send and receive. The same PetscSF can be applied to different root/leaf data or data types. Depending on the communication pattern, PetscSF can perform the operation using MPI point to point or collective operations. For the most common sparse neighborhood communication pattern, we use persistent MPI sends and receives by default, but we also support MPI neighborhood collectives.
@@ 4.2 The mismatch between MPI and GPUs
As more and more computations are offloaded to device for acceleration, it is desirable to directly communicate data on device instead of staging data on host just for the purpose of communication, because moving data between host and device is expensive. GPU-aware MPI was introduced to solve this problem by allowing MPI calls to accept device buffers. However, one has to be aware that GPU computations or kernels are queued in streams and are executed asynchronously with respect to the host process. MPI routines, as a host side API, must make sure the send data on device, which could be the output of some previous GPU kernels, is ready to be sent. We would like to inform MPI of the stream in use so that it can maintain the data dependency. Unfortunately, as of this writing, the MPI standard does not support this functionality, since MPI routines do not take a GPU stream argument. As a result, MPI users must synchronize the GPU stream to have the send data ready: In other words, force all kernels producing the send data to complete before launching new kernels. On the receiver side, an MPI receive must block subsequent GPU kernel launches, i.e., synchronize the device. In PETSc, we have a default device stream. With that, Figure 2 gives a typical data path when using CUDA for PetscSFReduce(), which reduces sparse leaf data to root data with the help of the Pack/Unpack kernels.
@@
Because kernel launches are expensive, the GPU runtime provides users the stream mechanism for pipelining kernel launches, so that kernel launch latency can be hidden by the execution of previous kernels. MPI incurs a device synchronization that stalls the pipeline, making the cost of kernel launches harder to mitigate. To quantify the launch latency, we designed a microbenchmark, in which we launch an empty GPU kernel many times, marking on the CPU before and after each launch, and then calculate the average latency per launch (i.e., per iteration). We had two variants: in one (Asynchronous), there were no stream synchronizations at all; in the other (Synchronous), we inserted a stream synchronization after each launch. The results are shown in Table 2, from which we can see the huge impact of synchronization on kernel launch latency.
@@ 4.3 Synchronization-free PetscSF with NVSHMEM
To avoid the device synchronizations imposed by MPI, we developed an experimental implementation of PetscSF using NVSHMEM (NVIDIA, 2024). NVSHMEM is NVIDIA’s implementation of OpenSHMEM (Open Source Software Solutions, Inc., 2020) on CUDA devices. The OpenSHMEM standard specifies an API for partitioned global address space (PGAS) parallel programming, providing one-sided shared-memory style put/get APIs to access remote objects. NVSHMEM supports point-to-point and collective communications between GPUs within a node or over networks. Communication can be initiated either on the host or on the device. Unlike MPI, NVSHMEM host-side APIs take a stream argument. Remotely accessible objects (aka symmetric objects) are collectively allocated over all processing elements (PEs, like processes in MPI) from a special heap (called the symmetric heap). All PEs must allocate a symmetric object with the same size, so that the object always appears at the same offset in their symmetric heap. PEs access remote data by referencing a symmetric address and the rank of the remote PE. A symmetric address is the address of a symmetric object on the local PE, plus an offset if needed. In the NVSHMEM implementation of PetscSF, we still pack/unpack data in send/receive buffers, but those buffers are now allocated as symmetric objects on device. We launch, on the PETSc default stream, a CUDA kernel which then calls the device function nvshmem_putmem_nbi to put data from send buffers into receive buffers. NVSHMEM can be used with MPI. In our design we use MPI to communicate information, such as sizes and offsets of buffers, to help set up auxiliary data structures. We use NVSHMEM only in real communication. We found implementing PetscSF in NVSHMEM no less complex than in MPI. Symmetric allocation forced us to take the maximal size of send/receive buffers across all PEs, though their real sizes could vary. Shared-memory style APIs forced us to design complex communication protocols even for simple patterns, such as knowing when data in a send buffer has already been sent and thus is safe to overwrite, or when data in a receive buffer is ready for use. The detailed design was published in Zhang et al. (2021). We hope in the near future that the MPI Forum will provide GPU-initiated communication support in the MPI standard, such as in the proposal of Zhou et al. (2022), which would greatly simplify implementation. Currently, our PetscSF implementation is capable of doing stream-aware communication across CUDA devices, free of device synchronization. In Section 6.1, we have used it to implement a parallel asynchronous conjugate gradient (CG) solver.
4.4 GPU message passing latency on (pre-)exascale machines
@@
To measure GPU-aware MPI communication performance on various machines, we designed a microbenchmark shown in Listing 4. One of the star-forests used is illustrated in the right of Figure 1, where there are n consecutive roots on the first process and n consecutive leaves on the second. The leaves are connected to the roots one-on-one in order. The code works as if the two processes keep bouncing a message of n*sizeof(double) bytes to the other.
@@
When op is MPI_REPLACE, the two buffers, rdata for root data and ldata for leaf data, are directly used as MPI send/receive buffers. The communication in this case is very similar to the osu_latency test from the OSU Micro-Benchmarks (Panda et al., 2024), with the exception that we synchronize the device before sending messages for reasons mentioned in Section 4.2, while osu_latency does not. We named the test SF-pingpong and measured the one-way latency of a message, with results on the four (pre-)exascale machines shown in Figure 3. Note that on Frontier and Aurora, we treated the two GCDs (or tiles) within a GPU package as two separate GPUs. In all tests, we used the two closest GPUs within a compute node, in other words, the two are either in the same package or connected to the same CPU socket (as on Summit or Perlmutter). For comparison, we also give the intra-node CPU to CPU message latency on Perlmutter measured with the same code with host data. From Figure 3 we can see MPI latency with device data is generally much higher than that with host data for small and medium messages, which implies GPU operations have a high start-up cost. We can see that Perlmutter has much lower latency than Summit thanks to the hardware and software upgrade listed in Table 1.
Figure 3: SF-pingpong test: MPI latency between two closest GPUs on the four (pre-)exascale machines, and between two CPU cores within a compute node on Perlmutter. In the legend parentheses are GPU vendor names. Note the strikingly better performance on Frontier with small messages, compared with other machines.
@@ 5 Portable matrix assembly on GPUs
In Section 3, we previewed some new PETSc matrix assembly APIs for GPUs in the Jacobian computation. In this section we discuss them in detail. PETSc has a rich set of APIs for CPU (host) matrix assembly. In the past, to achieve best performance users needed to call MatXAIJSetPreallocation() in advance to preallocate memory for the matrix, but the introduction of a hash table-based matrix assembly approach has rendered explicit preallocation no longer necessary in most cases. Next, users call MatSetValues() to insert blocks of values, e.g. element matrices from finite elements, into the matrix using global indices. These values could be local (i.e., owned by the calling MPI process) or remote (i.e., owned by other processes), determined by the layout of the matrix. For local values, PETSc directly inserts them, while for remote ones, PETSc might stash them until MatAssemblyBegin()/End() are called, when MPI communication is used to distribute stashed values and the matrix is finally assembled. While these APIs are convenient, they are all host APIs working with host memory, and can be called only by a single CPU thread. Directly porting MatSetValues() and related functions to a GPU is not feasible, since if multiple GPU threads call MatSetValues(), PETSc would need to stash remote values (and potentially allocate new stash area on device), do binary searches to find locations to insert local values, and use atomics to prevent data races when multiple GPU threads insert to the same non-zero location. Though these operations keep memory utilization low, they are too latency-sensitive for efficient computing on GPUs. Various approaches, like coloring to avoid atomics or using lookup tables to avoid binary search, were proposed (Cecka et al., 2011; Trotter et al., 2023). But these were for a single process and in effect shifted the burden to application developers. To support MPI parallelism and provide users convenience close to MatSetValues(), we designed a new set of coordinate-based (COO) matrix assembly APIs for matrices in the popular compressed sparse row (CSR) format. In both PETSc’s native and GPU formats, as well as Hypre’s ParCSR, MPI parallel matrices are distributed row-wise across MPI processes with diagonal (intra-process coupling) and off-diagonal (inter-process coupling) blocks stored separately as two sequential CSR matrices. The classic COO format consists of three arrays, i[], j[], v[] of equal length, in which the assembled matrix $A$ is defined as the sum of each contribution v[k] to entry $a_{\texttt{i[k]},\texttt{j[k]}}$, so that each index $k$ represents a distinct nonzero. In nonlinear and transient solves, one needs to repeatedly assemble matrices with the same nonzero pattern but different numeric values.
@@
PETSc splits COO assembly into a symbolic preallocation stage with MatSetPreallocationCOO(A,n,i,j), and one or more numeric stages with MatSetValuesCOO (A,v,mode). During preallocation, it analyzes indices in i/j[] of length n on the host, exchanges information about remote entries, finalizes the sparsity pattern of the diagonal and off-diagonal blocks, preallocates device memory, and builds MPI communication plans. The arrays i/j[] can be freed after this stage. We allow negative indices in i/j[], meaning the corresponding entries will be ignored, a convenient way to handle boundary conditions. MatSetValuesCOO() sets elements of the matrix, where v[] is an array on device that has the same length and follows the same order as i/j[]. Each entry (with non-negative indices) is destined for the owned diagonal, owned off-diagonal block, or a send buffer. The implementation first calls a kernel to fill the send buffer on device and initiates the GPU-aware MPI communication, then calls two asynchronous kernels filling nonzeros in the diagonal and off-diagonal blocks, in which each thread accumulates into a single nonzero entry. After completing communication, the implementation calls two similar kernels unpacking entries from the receive buffer. With knowledge of the sparsity pattern and insertion order in advance, our implementation avoids data races and atomics completely with some helper data structures.
@@
With PETSc’s COO matrix assembly, users need to provide i[], j[] and v[] at once. In finite element assembly, they can loop over elements and precompute i/j[] on host, assuming entries in the same element matrix will be stored contiguously in the arrays. Meanwhile, to prepare for the concurrent computing of element matrices on device, one also needs to know the offset of each element in the array v[]. With homogeneous elements so that each element matrix is the same size, the offsets can be computed analytically. Otherwise, one can create an auxiliary array, say offset[], to store the information and also copy it to the device. Then in the Jacobian, each GPU thread in charge of one element conveniently gets a pointer by &v[offset[tid]], to store element matrix values, where tid is the thread id.
@@
The above COO matrix assembly APIs are portable across host and device, so that the value array v[] can be on device or on the host, depending on the type of the matrix. For example, if $A$’s type is MATAIJ, a PETSc host CSR matrix type, then v[] needs to be on host. If it is MATAIJCUSPARSE, a PETSc matrix type for CUDA devices, then v[] needs to be in CUDA device memory. If it is MATAIJKOKKOS, a PETSc matrix type provided by the PETSc/Kokkos backend, then v[] needs to be in the Kokkos default memory space. To be complete, we also provide VecSetPreallocationCOO(x,n,i) and VecSetValuesCOO(x,v,mode) for COO vector assembly on device.
6 Asynchronous solvers on GPUs
@@ 6.1 Distributed asynchronous CG with PetscSF over NVSHMEM
6.2 Improve asynchronicity with PetscDeviceContext and managed memory

### v2

@@ 4 Communication on GPUs
As PETSc’s strength lies in distributed computation with sparse and irregular data, the library has to take care of complex communication among processes, such as those in sparse matrix-vector multiplication, sparse matrix-matrix multiplication, or irregular mesh partitioning. In PETSc, we designed a module named PetscSF (Zhang et al., 2021) to encapsulate frequently used communication operations. Underneath a unified interface, PetscSF can have different communication implementations, though the default uses MPI. In this section, we first introduce PetscSF’s design and the extensions to allow MPI communication with device data. Next, we introduce an experimental PetscSF implementation using NVIDIA NVSHMEM (NVIDIA, 2024) that can overcome some limitations imposed by MPI. Finally, we use a microbenchmark written with PetscSF to study the GPU-aware MPI communication performance on the four target machines.
@@ 4.1 The star-forest (SF) abstraction
PetscSF uses star-forests to abstract communication patterns. A star is a simple tree with one root vertex connected to zero or more leaves. A star forest is a disjoint union of stars; see examples in Figure 1.
@@
The former broadcasts root values to leaves, while the latter reduces leaf values into roots, with both taking a reduction operation specified by an MPI_Op argument to add the source values or to replace the destination values (with op=MPI_REPLACE). The Begin/End split-phase design allows users to insert computations in between to potentially overlap computation with communication. It is common in irregular applications that leaves and roots are not consecutive in their respective index space. In that case, PETSc will call its pack or unpack kernels to put data into internally managed buffers for send and receive. The same PetscSF can be applied to different root/leaf data or data types. Depending on the communication pattern, PetscSF can perform the operation using MPI point-to-point or collective operations. For the most common sparse neighborhood communication pattern, we use persistent MPI sends and receives by default, but we also support MPI neighborhood collectives.
@@ 4.2 The mismatch between MPI and GPUs
As more and more computations are offloaded to the device for acceleration, it is desirable to directly communicate data on the device instead of staging data on the host merely for the purpose of communication, because moving data between host and device is expensive. GPU-aware MPI was introduced to solve this problem by allowing MPI calls to accept device buffers. However, one has to be aware that GPU computations or kernels are queued in streams and are executed asynchronously with respect to the host process. MPI routines, as a host side API, must make sure that the send data on the device, which could be the output of some previous GPU kernels, is ready to be sent. We would like to inform MPI of the stream in use so that it can maintain the data dependency. Unfortunately, as of this writing, the MPI standard does not support this functionality since MPI routines do not take a GPU stream argument. As a result, MPI users must synchronize the GPU stream to have the send data ready: In other words, force all kernels producing the send data to complete before launching new kernels. On the receiver side, an MPI receive must block subsequent GPU kernel launches, i.e., synchronize the device. In PETSc, we have a default device stream. With that, Figure 2 gives a typical data path when using CUDA for PetscSFReduce(), which reduces sparse leaf data to root data with the help of the Pack/Unpack kernels.
@@
Because kernel launches are expensive, the GPU runtime provides users the stream mechanism for pipelining kernel launches, so that kernel launch latency can be hidden by the execution of previous kernels. MPI incurs a device synchronization that stalls the pipeline, making the cost of kernel launches harder to mitigate. To quantify the launch latency, we designed a microbenchmark, in which we launch an empty GPU kernel many times, marking on the CPU before and after each launch, and then calculating the average latency per launch (i.e., per iteration). We had two variants: in one (Asynchronous), there were no stream synchronizations at all; in the other (Synchronous), we inserted a stream synchronization after each launch. The results are shown in Table 2, from which we can see the huge impact of synchronization on kernel launch latency.
@@ 4.3 Synchronization-free PetscSF with NVSHMEM
To avoid the device synchronizations imposed by MPI, we developed an experimental implementation of PetscSF using NVSHMEM (NVIDIA, 2024). NVSHMEM is NVIDIA’s implementation of OpenSHMEM (Open Source Software Solutions, Inc., 2020) on CUDA devices. The OpenSHMEM standard specifies an API for partitioned global address space (PGAS) parallel programming, providing one-sided shared-memory style put/get APIs to access remote objects. NVSHMEM supports point-to-point and collective communications between GPUs within a node or over networks. Communication can be initiated either on the host or on the device. Unlike MPI, NVSHMEM host-side APIs take a stream argument. Remotely accessible objects (aka symmetric objects) are collectively allocated over all processing elements (PEs, like processes in MPI) from a special heap (called the symmetric heap). All PEs must allocate a symmetric object with the same size, so that the object always appears at the same offset in their symmetric heap. PEs access remote data by referencing a symmetric address and the rank of the remote PE. A symmetric address is the address of a symmetric object on the local PE, plus an offset if needed. In the NVSHMEM implementation of PetscSF, we still pack/unpack data in send/receive buffers, but those buffers are now allocated as symmetric objects on the device. We launch, on the PETSc default stream, a CUDA kernel which then calls the device function nvshmem_putmem_nbi to put data from send buffers into receive buffers. NVSHMEM can be used with MPI. In our design, we use MPI to exchange information, such as sizes and offsets of buffers, required to set up the communication, and then we use NVSHMEM to carry it out. We found implementing PetscSF in NVSHMEM no less complex than in MPI. Symmetric allocation forced us to take the maximal size of send/receive buffers across all PEs, though their real sizes could vary. Shared-memory style APIs forced us to design complex communication protocols even for simple patterns, such as knowing when data in a send buffer has already been sent and thus is safe to overwrite, or when data in a receive buffer is ready for use. The detailed design was published in Zhang et al. (2021). We hope that the MPI Forum will soon provide GPU-initiated communication support in the MPI standard, such as in the proposal of Zhou et al. (2022), which would greatly simplify implementation. Currently, our PetscSF implementation can do stream-aware communication across CUDA devices, free of device synchronization. In Section 6.1, we have used it to implement a parallel asynchronous conjugate gradient (CG) solver.
4.4 GPU-aware MPI message passing latency on (pre-)exascale machines
@@
Although the experiments using NVSHMEM are promising, GPU-aware MPI is still PETSc’s default communication mechanism on the GPU, and we therefore designed the microbenchmark shown in Listing 4 to measure its performance. One of the star-forests used is illustrated in the right of Figure 1, where there are n consecutive roots on the first process and n consecutive leaves on the second. The leaves are connected to the roots one-on-one in order. The code works as if the two processes keep bouncing a message of n*sizeof(double) bytes to the other.
@@
When op is MPI_REPLACE, the two buffers, rdata for root data and ldata for leaf data, are directly used as MPI send/receive buffers. The communication in this case is very similar to the osu_latency test from the OSU Micro-Benchmarks (Panda et al., 2024), with the exception that we synchronize the device before sending messages for reasons mentioned in Section 4.2, while osu_latency does not. We named the test SF-pingpong and measured the one-way latency of a message, with results on the four (pre-)exascale machines shown in Figure 3. Note that on Frontier and Aurora, we treated the two GCDs (or tiles) within a GPU package as two separate GPUs. In all tests, we used the two closest GPUs within a compute node, in other words, the two are either in the same package or connected to the same CPU socket (as on Summit or Perlmutter). For comparison, we also give the intra-node CPU-to-CPU message latency on Perlmutter measured with the same code with host data. From Figure 3, we can see that MPI latency with device data is generally much higher than that with host data for small and medium messages, which implies GPU operations have a high start-up cost. We can see that Perlmutter has much lower latency than Summit, presumably due to its newer hardware and software (see Table 1), though we do not know enough about the workings of IBM Spectrum MPI to say exactly why.
Figure 3: SF-pingpong test: MPI latency between two closest GPUs on the four (pre-)exascale machines, and between two CPU cores within a compute node on Perlmutter. Note the strikingly better performance on Frontier with small messages compared with other machines.
@@ 5 Portable matrix assembly on GPUs
In Section 3, we previewed some new PETSc matrix assembly APIs for GPUs in the Jacobian computation. In this section we discuss them in detail. PETSc has a rich set of APIs for CPU (host) matrix assembly. In the past, to achieve best performance users needed to call MatXAIJSetPreallocation() in advance to preallocate memory for the matrix, but the introduction of a hash table-based matrix assembly approach has rendered explicit preallocation no longer necessary in most cases. Next, users call MatSetValues() to insert blocks of values, e.g. element matrices from finite elements, into the matrix using global indices. These values could be local (i.e., owned by the calling MPI process) or remote (i.e., owned by other processes), determined by the layout of the matrix. For local values, PETSc directly inserts them, while for remote ones, PETSc might stash them until MatAssemblyBegin()/End() are called, when MPI communication is used to distribute stashed values and the matrix is finally assembled. While these APIs are convenient, they are all host APIs working with host memory, and can be called only by a single CPU thread. Directly porting MatSetValues() and related functions to a GPU is not feasible, since if multiple GPU threads call MatSetValues(), PETSc would need to stash remote values (and potentially allocate new stash area on device), do binary searches to find locations to insert local values, and use atomics to prevent data races when multiple GPU threads insert to the same non-zero location. Though these operations keep memory utilization low, they are too latency-sensitive for efficient computing on GPUs. Various approaches, like coloring to avoid atomics or using lookup tables to avoid binary search, were proposed (Cecka et al., 2011; Trotter et al., 2023). But these were for a single process and, in effect, shifted the burden to application developers. To support MPI parallelism and provide users convenience close to MatSetValues(), we designed a new set of coordinate-based (COO) matrix assembly APIs for matrices in the popular compressed sparse row (CSR) format. In both PETSc’s native and GPU formats, as well as Hypre’s ParCSR, MPI parallel matrices are distributed row-wise across MPI processes with diagonal (intra-process coupling) and off-diagonal (inter-process coupling) blocks stored separately as two sequential CSR matrices. The classic COO format consists of three arrays, i[], j[], v[] of equal length, in which the assembled matrix $A$ is defined as the sum of each contribution v[k] to entry $a_{\texttt{i[k]},\texttt{j[k]}}$, so that each index $k$ represents a distinct nonzero. In nonlinear and transient solves, one needs to repeatedly assemble matrices with the same nonzero pattern but different numeric values.
@@
PETSc splits COO assembly into a symbolic preallocation stage with MatSetPreallocationCOO(A,n,i,j), and one or more numeric stages with MatSetValuesCOO (A,v,mode). During preallocation, it analyzes indices in i/j[] of length n on the host, exchanges information about remote entries, finalizes the sparsity pattern of the diagonal and off-diagonal blocks, preallocates device memory, and builds MPI communication plans. The arrays i/j[] can be freed after this stage. We allow negative indices in i/j[], meaning the corresponding entries will be ignored, a convenient way to handle boundary conditions. MatSetValuesCOO() sets elements of the matrix, where v[] is an array on the device that has the same length and follows the same order as i/j[]. Each entry (with non-negative indices) is destined for the owned diagonal, owned off-diagonal block, or a send buffer. The implementation first calls a kernel to fill the send buffer on the device and initiates the GPU-aware MPI communication, then calls two asynchronous kernels filling nonzeros in the diagonal and off-diagonal blocks, in which each thread accumulates into a single nonzero entry. After completing communication, the implementation calls two similar kernels unpacking entries from the receive buffer. With knowledge of the sparsity pattern and insertion order in advance, our implementation completely avoids data races and atomics with some helper data structures.
@@
With PETSc’s COO matrix assembly, users need to provide i[] and j[] at once. In finite element assembly, they can loop over elements and precompute i/j[] on the host, assuming entries in the same element matrix will be stored contiguously in the arrays. Meanwhile, to prepare for the concurrent computing of element matrices on the device, one also needs to know the offset of each element in the array v[]. With homogeneous elements so that each element matrix is the same size, the offsets can be computed analytically. Otherwise, one can create an auxiliary array, say offset[], to store the information and also copy it to the device. Then in the Jacobian, each GPU thread in charge of one element conveniently gets a pointer by &v[offset[tid]], to store element matrix values, where tid is the thread id.
@@
The above COO matrix assembly APIs are portable across host and device, so that the value array v[] can be on device or on the host, depending on the type of the matrix. For example, if $A$’s type is MATAIJ, a PETSc host CSR matrix type, then v[] needs to be on the host. If it is MATAIJCUSPARSE, a PETSc matrix type for CUDA devices, then v[] can be host or device memory. If it is MATAIJKOKKOS, a PETSc matrix type provided by the PETSc/Kokkos backend, then v[] needs to be in the Kokkos default memory space. To be complete, we also provide VecSetPreallocationCOO(x,n,i) and VecSetValuesCOO(x,v,mode) for COO vector assembly on device.
6 Asynchronous linear solvers on GPUs
@@ 6.1 Distributed asynchronous CG with PetscSF over NVSHMEM
6.2 Improve solver asynchronicity with PetscDeviceContext and managed memory